Calcule ΔUrxn (kJ/mol) a 298,15 K para la siguiente reacción:

CH3OH (g) + CO (g) → CH3COOH (l)


1
Expert's answer
2021-03-11T06:55:45-0500

The change in enthalpy upon reaction of 51.90 g of Cu2O(s) is –52.96 kJ. Calculate the work, w, and energy ... "n(CH3OH) = m(CH3OH)\/M(CH3OH)" =51.9/144 = 0.36 mol ... ΔUrxn = –52.51 kJ.